# On the Surprising Effectiveness of Spectral Clipping in Learning Stable Linear Dynamics

### v4

@@ Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to enforce stability often preserve accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix that was computed using unconstrained least squares. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping any eigenvalues that are larger than one to one (without altering the eigenvectors). Through comprehensive experiments involving two different applications and publicly available benchmark datasets, we show that this simple technique can efficiently learn highly-accurate linear systems that are provably-stable. Notably, we find that SC can match or outperform strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Finally, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations.
Keywords: Linear Dynamical Systems, Stability, System Identification
@@ 1 Introduction
In spite of the growing dominance of deep neural networks, modeling and learning of linear dynamical systems (LDS) remain relevant due to the availability of analytical solutions, significantly lower computational burden, and ease of analyses and inspection. LDS continue to find use in diverse applications such as computer vision [1, 2, 3, 4] and time series prediction [5, 6]. Even when the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory [7, 8, 9] to approximately represent nonlinear systems as linear systems in a higher-dimensional lifted state space. These techniques have further increased the relevance and importance of learning LDS in complex applications, such as legged robotics [10, 11], robotic manipulation [12, 13, 14, 15], atmospheric sciences [16, 17, 18], and fluid dynamics [19, 20].
@@
When time-series data associated with system evolution is available, learning LDS can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective [21]. However, such unconstrained optimization methods tend to ignore important properties of the underlying systems, such as stability. Consequently, even if the underlying system is known to be stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
@@
A number of methods have been developed to enforce stability of learned LDS and reduce sensitivity to training data [22, 23, 2, 24, 25, 26, 27]. These methods take one of three approaches: a) minimizing the regression objective under stability constraints [22, 23], b) iterating between solving an unconstrained problem and refining the solution under stability constraints [2, 24], and c) iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective [25, 26, 27]. Regardless of the underlying approach, these methods inevitably introduce a considerable computational burden, potentially countering one of the primary advantages of leveraging linear techniques.
@@
In this work, we investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn stable LDS from data while introducing negligible additional computations. SC involves a straightforward implementation: 1) compute the system matrix via unconstrained least-squares, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. Our investigation is inspired by prior works that have revealed the surprising effectiveness of linear techniques: controlling the singular values of linear layers improves training stability and generalization error [28, 29, 30]; retraining only the last linear layer improves robustness to spurious correlations [31]; and Low-Rank Adaptation (LoRA) dramatically improves efficiency [32].
@@
SC enforces stability by construction without requiring constrained optimization or iterative updates. In contrast to existing approaches discussed above, SC can efficiently scale to high-dimensional systems (e.g., image prediction). Further, seminal works on the Koopman operator [33, 19, 9] reveal fundamental connections between the spectrum of the Koopman matrix and desirable properties in learned nonlinear systems. SC can be trivially combined with Koopman operators to enhance the predictive accuracy and stability of learned nonlinear dynamical systems (e.g., robotic systems).
@@
We conducted comprehensive experiments involving benchmark datasets across different domains (two video prediction datasets [1, 34] and one robot model learning dataset [35]) and strong baselines (CG [2], WLS [24], and SOC [26]). Our results reveal that SC consistently achieves the lowest reconstruction error, and generates stable and more accurate long-term predictions. Notably, SC is able to achieve this level of performance while being orders-of-magnitude faster.
@@
Additional experiments show that SC can be readily combined with Koopman operators to learn stable nonlinear systems (such as those that underlie complex dexterous manipulation skills for a multi-fingered robotic hand [13]) in mere seconds. Importantly, we show that SC enables the learned system to remain robustly stable and effective even the training data includes unsuccessful demonstrations or when the task horizon is truncated (characteristics are commonly observed when robot learning data is collected from humans [36, 37, 38]). In contrast, we find that unconstrained learning produces unstable LDS that lead to erratic and ineffective robot behaviors when such challenging characteristics are present in the training data.
@@
In summary, our core contributions include: i) a simple post-hoc approach to learn stable linear dynamical systems, ii) evidence for its surprising ability to simultaneously offer reliable long-term predictions, high computational efficiency, and stability guarantees, iii) the ability to trade off between expressivity and stability, iv) an extension leveraging Koopman operators to accommodate stable nonlinear systems, and v) robustness against unsuccessful and truncated demonstrations.
2 Problem Formulation
Consider a stable discrete-time autonomous linear dynamical system $\mathrm{x}_{t+1}=A\mathrm{x}_{t}$, where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state at time $t$, and $A\in\mathcal{S}_{A}$ is the unknown system matrix, and $\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the subspace that contains stable matrices: $\mathcal{S}_{A}\triangleq\{A\in\mathbb{R}^{n\times n}|\max\{|\lambda_{i}(A)|\}^{n}_{i=1}<=1\}$, where $\{\lambda_{i}(A)\}^{n}_{i=1}$ are the eigenvalues of the matrix $A$, and $|\lambda|$ is the $|\lambda|\coloneqq\sqrt{\text{Re}(\lambda)^{2}+\text{Im}(\lambda)^{2}}$.
Unconstrained Learning: Suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T}]$ containing the time evolutions of the above stable LDS. Defining $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, allows one to express the problem of learning the above LDS from a dataset $D$ as the following solving standard least-square regression problem [39, 40]:
$\hat{A}_{LS}=\arg\min_{\Theta}||Y-\Theta X||_{F}^{2},$ (1)
@@ 2 Problem Formulation
where $\Theta\in\mathbb{R}^{n\times n}$ are the learnable parameters and $||\cdot||_{F}$ is the matrix Frobenius norm. However, optimizing the Eq. (1) ignores the constraints on the matrix $A$, such as the system stability that is crucial for the accurate long-term prediction.
@@
Learning Stable LDS: Our goal is to ensure that the learned system respects the stability constraints of the underlying unknown LDS, despite learning from the limited dataset $D$. To this end, we can rewrite the optimization problem Eq. (1) to include stability constraints:
$\hat{A}_{S}=\arg\min_{\Theta\in\mathcal{S}_{A}}||Y-\Theta X||_{F}^{2},$ (2)
@@
While solving the optimization problem in Eq. (2) can yield the constrained optimal solution which is guaranteed to be stable, it is known to be highly non-convex and requires approximations [41, 26].
3 Spectrum Clipping: Enforcing Stability of Learned LDS
We evaluate a simple method, which we refer to as Spectrum Clipping (SC), to ensure the stability of learned LDS without having to directly solve the constrained optimization in Eq. (2).
@@ 3 Spectrum Clipping: Enforcing Stability of Learned LDS
Eigen Decomposition of Unconstrained Solution: First, we compute the system evolution matrix $\hat{A}_{LS}$ by optimizing the least square regression (Eq. (1)). Second, we perform eigen decomposition of the least squares solution: $\hat{A}_{LS}=M\Lambda M^{-1}$, where $\Lambda$ is a diagonal matrix with all the eigenvalues of $\hat{A}_{LS}$ on its diagonal, and $M$ is the modal matrix with the corresponding eigenvectors.
Spectral Clipping: Next, we identify all unstable eigenvalues in $\Lambda$ and clip their magnitudes to $1$: $\bar{\Lambda}_{i}=1,\forall i=j,\text{such that}|\Lambda_{j}|>1$ , and leave all other eigenvalues in $\Lambda$ unaltered: $\bar{\Lambda}_{i}=\Lambda_{i},\forall i=j,\text{such that}|\Lambda_{j}|\leq 1$. Finally, we reconstruct the system matrix based on the new $\bar{\Lambda}$ and the unaltered modal matrix $M$ to produce SC’s estimate of the system matrix: $\hat{A}_{SC}\triangleq M\bar{\Lambda}M^{-1}$.
Note that the above post-hoc clipping procedure enforces stability constraints by ensuring that $\hat{A}_{SC}\in\mathcal{S}_{A}$ without requiring constrained optimization. Indeed, it is natural to suspect that this seemingly-naive approach would result in significantly poorer, albeit stable, approximations and worse prediction accuracies. However, as our experiments demonstrate, SC surprisingly leads to similar, if not better, accuracy compared to existing constrained optimization techniques.
Ensuring Lyapunov Stability: Enforcing Since SC clips eigenvalues to 1, the resulting system may not be strictly stable in the sense of Lyapunov. If desired, SC can be trivially extended to ensure Lyapunov stability by merely clipping the magnitude of the eigenvalues to marginally less than $1$. We formalize this by clipping the magnitudes to $1-\varepsilon$ (instead of $1$), where $\varepsilon\in[0,1]$ and Lyapunov stability is ensured when $\varepsilon>0$.
Expressivity-Stability Tradeoff: We note that the choice of $\varepsilon$ above represents an important tradeoff between the system’s expressivity and stability. In particular, choosing $\varepsilon=0$ will yield marginally-stable but highly-expressive systems. In contrast, choosing $\varepsilon>0$ will ensure strictly-stable system, with increasing values of $\varepsilon$ yielding faster exponential convergence and diminishing expressivity (see Appendix. E for experimental validation).
Dealing with Non-diagonalizable Unconstrained Solutions: A potential concern about the above approach is that $\hat{A}_{LS}$ could be non-diagonalizable. However, it is well known that one can always first find a diagonalizable matrix arbitrarily close to the original non-diagonalizable matrix (see Lemma. 3.1).
Lemma 3.1.
For any $\gamma>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. $||\hat{A}_{LS}-\tilde{A}_{LS}||_{1}<\gamma$.
Proof.
Lemma 3.1 can be proved by the fact that the set of real matrices that are diagonalizable over $\mathbb{C}$ is dense in $\mathbb{R}^{n\times n}$. The details can be found from the matrix perturbation theory [42, 43]. ∎
@@
The above result can be trivially extended to show that the evolution of the system with a non-diagonalizable matrix and that of the system with the arbitrarily-close diagonalizable matrix can be made arbitrarily close (see proof in Appendix. A). As such, the SC method can be applied to the arbitrarily-close diagonalizable matrix $\tilde{A}_{LS}$ to enforce stability of systems with non-diagonalizable matrices without significant approximation errors.

### v6

@@ Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) verifiable stability, and iii) computational efficiency. Unconstrained minimization of prediction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to enforce stability often preserve accuracy, but do so only at the cost of increased computation. In this work, we investigate if a seemingly-naive procedure can simultaneously offer all three desiderata. Specifically, we consider a post-hoc procedure in which we surgically manipulate the spectrum of the linear system after it was learned using unconstrained least squares. We call this approach spectral clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after any eigenvalues whose magnitude exceeds one have been clipped to one (without altering the eigenvectors). We also show that SC can be readily combined with Koopman operators to learn nonlinear dynamical systems that can generate stable predictions of nonlinear phenomena, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Through comprehensive experiments involving two different applications and publicly available benchmark datasets, we show that this simple technique can efficiently learn highly-accurate predictive dynamics that are provably-stable. Notably, we find that SC can match or outperform strong baselines while being orders-of-magnitude faster. Finally, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations. Our code and datasets can be found at https://github.com/GT-STAR-Lab/spec_clip.
keywords:
Linear Dynamical Systems, Stability, System Identification.
@@ 1 Introduction
In spite of the growing dominance of deep neural networks, modeling and learning of linear dynamical systems (LDS) remain relevant due to the availability of analytical solutions, significantly lower computational burden, and ease of analyses and inspection. LDS continue to find use in diverse applications such as computer vision (Chan and Vasconcelos, 2005; Boots et al., 2007; Chien et al., 2022; Wang and Mu, 2024) and time series prediction (Liu et al., 2024; Takeishi et al., 2017). Even if the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory (Koopman, 1931; Williams et al., 2015; Mezic, 2020) to approximate nonlinear systems with linear systems in a higher-dimensional lifted state space. These techniques further increase the relevance and importance of learning LDS in complex applications, such as legged robotics (Kim et al., 2024; Li et al., 2024), robotic manipulation (Bruder et al., 2021; Han et al., 2023, 2024; Chen et al., 2024), atmospheric sciences (Navarra et al., 2021; Millard et al., 2024; Zheng et al., 2022), and fluid dynamics (Mezić, 2013; Schmid, 2022).
@@
When time-series data associated with system evolution is available, learning LDS can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective (Ljung, ). However, such unconstrained optimization methods tend to ignore important properties of the underlying systems, such as stability. Consequently, even if the underlying system is known to be stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
@@
A number of methods have been developed to enforce stability of learned LDS and reduce sensitivity to training data (Lacy and Bernstein, 2002, 2003; Boots et al., 2007; Huang et al., 2016; Gillis et al., 2020; Mamakoukas et al., 2020, 2023). These methods take one of three approaches: a) minimizing the regression objective under stability constraints (Lacy and Bernstein, 2002, 2003), b) iterating between solving an unconstrained problem and refining the solution under stability constraints (Boots et al., 2007; Huang et al., 2016), and c) iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (Gillis et al., 2020; Mamakoukas et al., 2020, 2023). In general, these methods inevitably introduce a considerable computational burden, potentially countering one of the primary advantages of linear techniques.
@@
In this work, we investigate a simple yet surprisingly effective method, Spectral Clipping (SC), to learn stable LDS from data with little extra computation. SC involves a straightforward implementation: 1) compute the system matrix via unconstrained least-squares, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip the magnitude of eigenvalues only when larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors.
@@
In contrast to existing approaches discussed above, SC offers stability-by-construction without requiring constrained optimization or iterative updates. SC can efficiently scale to high-dimensional systems (e.g., image prediction). Further, seminal works on the Koopman operator (Rowley et al., 2009; Mezić, 2013; Mezic, 2020) reveal fundamental connections between the spectrum of the Koopman matrix and desirable properties in learned nonlinear systems. We show that SC can be combined with Koopman operators to enhance the predictive accuracy and stability of learned nonlinear dynamical systems (e.g., robotic systems).
While it can be readily seen that SC ensures stability by directly manipulating the eigenvalues, it might seem puzzling that it does not hurt the model’s predictive accuracy. To explain SC’s surprising effectiveness, we provide practical intuitions and theoretical insights (e.g., see Fig. 1 for a geometric interpretation of SC). Indeed, prior works have identified several linear techniques that defy expectations: controlling the singular values of linear layers improves training stability and generalization (Boroojeny et al., 2024; Sedghi et al., 2018; Senderovich et al., 2022); retraining only the last linear layer improves robustness to spurious correlations (Kirichenko et al., 2022); and Low-Rank Adaptation (LoRA) dramatically improves efficiency (Hu et al., 2022).
@@
We conducted comprehensive experiments involving benchmark datasets across different domains (two video prediction datasets (Chan and Vasconcelos, 2005; Hadji and Wildes, 2018), and a robot model learning dataset (Gaz et al., 2019)) and strong baselines (CG (Boots et al., 2007), WLS (Huang et al., 2016), and SOC (Mamakoukas et al., 2020)). Our results reveal that SC consistently achieves comparable (if not better) predictive accuracy, and generates stable long-horizon predictions. Notably, SC is able to achieve this level of performance while being orders-of-magnitude faster.
@@
Our experiments also show that SC can be readily combined with Koopman operators to learn nonlinear systems capable of stable and consistent predictions in mere seconds (e.g., behavioral dynamics of dexterous manipulation skills for a multi-fingered robotic hand (Han et al., 2023)). Importantly, we show that SC enables the learned system to generate robust and stable predictions even when the training data includes unsuccessful or truncated demonstrations (commonly characteristics of real-world robot learning datasets (Qin et al., 2022; Shaw et al., 2023; Bahl et al., 2022)). In contrast, we find unconstrained learning yields unstable LDS leading to erratic and ineffective robot behaviors in the presence of such challenging characteristics.
@@
In summary, our core contributions include: i) a simple post-hoc procedure to ensure stability of learned linear dynamical systems, ii) evidence for its surprising ability to simultaneously offer reliable predictions, high computational efficiency, and stability guarantees, iii) practical and theoretical groundings of its underlying functionality tunable trade-off between expressivity and stability, iv) an extension leveraging Koopman operators to benefit nonlinear systems, and v) enabling robustness against unsuccessful and truncated demonstrations.
2 Background: Learning Stable Linear Dynamical Systems
[…]
Unconstrained Learning: In its simplest form, the problem of learning an LDS from $D$ reduces to solving a standard least-squares problem. (Montgomery et al., 2021; Seber and Lee, 2012):
[…]
@@ 2 Background: Learning Stable Linear Dynamical Systems
where $\Theta\in\mathbb{R}^{n\times n}$ are the learnable parameters and $||\cdot||_{F}$ is the matrix Frobenius norm. However, the optimization in Eq. (1) does not impose any constraints on the matrix $A$, resulting in the learning of potentially unstable systems despite our knowledge that the original system is stable.
@@
Stability-Constrained Learning: The prevailing strategy to ensure that the learned LDS is stable when learning from a dataset $D$ involves rewriting the optimization problem Eq. (1) to explicitly include stability constraints:
[…]
@@
While solving the optimization problem in Eq. (2) can yield the constrained optimal solution which is guaranteed to be stable, it is known to be non-convex and requires approximations (Boyd and Vandenberghe, 2004; Mamakoukas et al., 2020).
3 Post-Hoc Stability via Spectral Clipping
We investigate a simple post-hoc procedure, dubber “Spectral Clipping (SC)", that surgically manipulates the spectrum of the system matrix $\hat{A}_{LS}$ in order to learn a stable LDS. Our motivation is to study computationally-cheap alternatives to standard approaches that explicitly solve a complex constrained optimization problem, such as the one in Eq. (2).
Figure 1: Geometric intuition: Spectral clipping enables stable evolution by clipping the spectral radius of the system matrix without altering its eigenvectors.
Algorithm: Spectral Clipping (SC) involves the following steps:
Step 0. Unconstrained LS: Estimate $\hat{A}_{LS}$ as the system matrix obtained from data by analytically solving the unconstrained least-squares problem (Eq. (1)).
@@ 3 Post-Hoc Stability via Spectral Clipping
Step 1. Decomposition: Perform eigen decomposition of the least squares solution: $\hat{A}_{LS}=M\Lambda M^{-1}$, where $\Lambda$ is a diagonal matrix with all the eigenvalues of $\hat{A}_{LS}$ on its diagonal, and $M$ is the modal matrix with the corresponding eigenvectors.
Step 2. Spectral Clipping: Identify all unstable eigenvalues of $\hat{A}_{LS}$ in $\Lambda$ and clip their magnitudes:
$\bar{\Lambda}_{ii}=\left\{\begin{array}[]{ll}\frac{\Lambda_{ii}}{|\Lambda_{ii}|}(1-\varepsilon),&|\Lambda_{ii}|\geq 1\\ \Lambda_{ii},&|\Lambda_{ii}|<1\end{array}\right$., where $\varepsilon\geq 0$ is an arbitrarily-small constant.
[…]
Stability: The reconstructed linear system after the above clipping procedure, $x_{t+1}=\hat{A}_{SC}\ x_{t}$ can be trivially shown to be asymptotically (marginally) stable When $\varepsilon>0$ ($\varepsilon=0$) since the maximum eigenvalue of $\hat{A}$ is strictly less than (equal to) one by construction (Robinson, 2012).
@@
Dealing with non-diagonalizable Matrices: A potential concern with the above approach is that $\hat{A}_{LS}$ may be non-diagonalizable. In practice, we can add tiny perturbations to the original $\hat{A}_{LS}$ to ensure a valid eigen-decomposition, which the built-in solvers in NumPy and MATLAB also perform. This is justified by the fact that a diagonalizable matrix can always be found arbitrarily close to any non-diagonalizable one (see Lemma. A.1 in Appendix. A). We also leverage this fact to show that the evolution of the system with a non-diagonalizable matrix and that of the system with the arbitrarily-close diagonalizable matrix can be made arbitrarily close (see proof in Appendix. A). As such, the SC method can be applied to the arbitrarily-close diagonalizable matrix $\tilde{A}_{LS}$ to enforce stability of systems with non-diagonalizable matrices without significant approximation errors.
Prediction accuracy: Indeed, it is natural to suspect that this seemingly-naive approach would result in significantly poorer, albeit stable, approximations that produce worse predictions. However, as our experiments demonstrate, SC surprisingly leads to similar accuracy compared to existing constrained optimization techniques (see Sec. 5 for experiments across a wide range of domains).
Expressivity-Stability tradeoff: We note that the choice of $\varepsilon$ above represents an important tradeoff between the system’s expressivity and stability. In particular, choosing $\varepsilon=0$ will yield marginally-stable but highly-expressive systems. In contrast, choosing $\varepsilon>0$ will ensure strictly-stable systems, with increasing values of $\varepsilon$ yielding faster exponential convergence and diminishing expressivity (see Appendix. C for experimental validation).
Intuition: While SC’s effectiveness may seem puzzling at surface, a deeper analysis offers some intuition. By altering only the eigenvalues while preserving eigenvectors, SC essentially impacts the “speeds” or “rates” of the system’s dynamics modes without qualitatively altering them. Specifically, SC likely preserves the geometric structure of the state space as the invariant subspace spanned by the eigenvectors remains unchanged while modifying how trajectories behave within them. Since eigenvalues only determine whether trajectories grow, decay, or oscillate, altering their magnitudes shifts the stability margin without altering the system’s directions of approach to divergence. Further, SC intervenes only when necessary to ensure stability. By only altering eigenvalues that are larger than one, SC reduces the rate of energy growth along specific paths (dictated by corresponding eigenvectors), which would have pushed the system towards instability without clipping. See Fig. 1 for this geometric intuition.
Extensions: We also extend SC to settings beyond learning stable LDS. Specifically, we discuss both learning nonlinear systems (see Sec. 4) and systems with control inputs (see Appendix. B).
4 A Koopman-based Adaptation of SC to Nonlinear Systems
We also investigate the potential utility of SC when learning nonlinear dynamical systems (NLDS) by combining it Koopman operator-based techniques (Koopman, 1931; Mezic, 2020).
Let $\xi_{k+1}=f(\xi_{k})$ be an unknown discrete-time NLDS with $\xi_{k}\in\Xi\subseteq\mathbb{R}^{n},\forall k\in\mathbb{N}$ and $D_{f}=\{[\xi_{1}^{(j)},\xi_{2}^{(j)},\cdots,\xi_{T}^{(j)}]\}_{j=1}^{N_{f}}$ be a dataset containing $N_{f}$ trajectories of the NLDS. Koopman operator-based learning approaches learn to approximately encode the NLDS as $\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$, where $\hat{K}\in\mathbb{R}^{m\times m}$ is the learned finite-dimensional approximation of the Koopman operator $\mathrm{z}_{k}=\phi(\xi_{k})\in\mathbb{R}^{m}$ is the “lifted" latent state produced by the so-called “lifting" function $\phi:\Xi\rightarrow\mathcal{Z}$ with $m\gg n$, and $\mathcal{Z}$ is the approximated Koopman-invariant subspace in which the dynamics appear linear (Koopman, 1931; Mezic, 2020).
While it is possible to learn or design an encoder and a decoder to lift and reconstruct the original state, other formulations allow for trivial decoding and enable further analysis. In particular, we investigate state-inclusive latent space designs (Proctor et al., 2018; Han et al., 2023; Johnson et al., 2025) which append the original state to the nonlinear lifted embedding (i.e., $\mathrm{z_{t}}=[\phi(\xi)^{T},\xi^{T}]^{T}$ with $\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$). As such, a trivial linear map can recover the original states from the latent states.
Spectral clipping of the Koopman matrix: Given a learned lifted latent LDS ($\mathrm{z}_{k+1}=\hat{K}\ \mathrm{z}_{k}$), we can perform spectral clipping of the learned Koopman matrix $\hat{K}$ following the same procedure described in Sec. 3. This process will yield a Schur-stable Koopman matrix $\hat{K}_{SC}$ with eigenvalues whose magnitudes are guaranteed to lie within the unit circle.
Open-loop predictive stability: Indeed, stability in the latent space does not necessarily imply stability in the original state space of the NLDS, and in fact requires assumptions about forward and backward conjugacy Lutkus et al. (2025). As such, we do not claim that SC ensures closed-loop stability of the underlying NLDS. However, SC ensures that lifted latent LDS is stable by construction in $\mathcal{Z}$. Given any initial condition ($\mathrm{z_{0}}=[\phi(\xi_{0})^{T},\xi_{0}^{T}]^{T}$), the predictions generated by the latent linear system is guaranteed to converge to the origin ($\hat{\mathrm{z}}_{k}\rightarrow 0$ as $k\rightarrow\infty$). Since we employ state-inclusive lifting ($\mathrm{\hat{z}_{k}}=[\phi(\hat{\xi}_{k})^{T},\hat{\xi}_{k}^{T}]^{T}$), the predicted original states $\hat{\xi}$ are merely a subset of the predicted latent states $\hat{\mathrm{z}}_{k}$ and as such will also converge to the origin ($\hat{\xi}_{k}\rightarrow 0$ as $k\rightarrow\infty$). We refer to this property as open-loop predictive stability.